# Enabling Vibration-Based Gesture Recognition on Everyday Furniture via Energy-Efficient FPGA Implementation of 1D Convolutional Networks

### v1

@@ IV-C Experiment 2: Hardware-Aware Optimization
In addition, prior work [9] reports that training their 2D-CNN model on six sessions of vibration data took approximately 7,611 seconds on a CPU. In contrast, each trial in our hardware-aware search (up to 100 epochs) completes in an average of 479.89 seconds for 1D-CNN and 472.99 seconds for 1D-SepCNN on an RTX 2080 SUPER GPU, demonstrating at least an order-of-magnitude improvement in training efficiency per deployable configuration. Regarding inference time, Yoshida reports a total latency of 365 ms, comprising 15 ms for STFT-based feature extraction and 350 ms for 2D-CNN inference. Our FPGA-deployed models achieve significantly lower inference latency. For instance, as low as 6.83 ms, representing a speedup of more than 53 over Yoshida’s. This drastic improvement in latency is achieved alongside substantial reductions in model size and power consumption, highlighting the real-time responsiveness and deployability of our embedded solution.
@@ V Conclusion and Future Work
This work presents a fully FPGA-deployable vibration-based gesture recognition solution for smart surfaces, targeting low-latency, energy-efficient interaction without reliance on costly preprocessing or high-performance CPUs with high energy usage. To this end, we propose a series of end-to-end software-hardware co-optimizations. We firstly eliminate complex spectral transformations by operating directly on downsampled raw vibration signals, reducing input size by 21 while preserving gesture-discriminative patterns. We then design two lightweight 1D convolutional architectures (1D-CNN and 1D-SepCNN) that dramatically reduce parameter counts (down to 216) compared to the 2D-CNN baseline (369 million), enabling compact yet expressive inference on resource-limited hardware. We extend the ElasticAI.Creator library to support these models for integer-only quantization and template-based RTL design generation. In addition, we integrate a hardware-aware Optuna framework to enable constraint-driven selection of model configurations that satisfy bi-objective deployment goals. Experimental results confirm that all selected configurations run under 21 ms and consume less than 3.2 mJ on the AMD Spartan-7 XC7S25 FPGA. The most efficient setting (1D-SepCNN, PS split) completes inference in under 7 ms with energy consumption smaller than 1.2 mJ, offering over 53 speedup compared to the CPU baseline [9]. While 1D-CNN generalizes better across users and tables, 1D-SepCNN delivers superior latency and energy efficiency, highlighting a trade-off between robustness and deployment cost. Together, these results validate the feasibility of real-time gesture recognition on low-power FPGA platforms.

### v2

@@ IV-C Experiment 2: Hardware-Aware Optimization
In addition, prior work [yoshida2023smatable2] reports that training their 2D-CNN model on six sessions of vibration data took approximately 7,611 seconds on a CPU. In contrast, each trial in our hardware-aware search (up to 100 epochs) completes in an average of 479.89 seconds for 1D-CNN and 472.99 seconds for 1D-SepCNN on an RTX 2080 SUPER GPU, demonstrating at least an order-of-magnitude improvement in training efficiency per deployable configuration. Regarding inference time, Yoshida reports a total latency of 365 ms, comprising 15 ms for STFT-based feature extraction and 350 ms for 2D-CNN inference. Our FPGA-deployed models achieve significantly lower inference latency. For instance, as low as 6.83 ms, representing a speedup of more than 53 over Yoshida’s. This drastic improvement in latency is achieved alongside substantial reductions in model size and power consumption, highlighting the real-time responsiveness and deployability of our embedded solution.
@@ V Conclusion and Future Work
This work presents a fully FPGA-deployable vibration-based gesture recognition solution for smart surfaces, targeting low-latency, energy-efficient interaction without reliance on costly preprocessing or high-performance CPUs with high energy usage. To this end, we propose a series of end-to-end software-hardware co-optimizations. We firstly eliminate complex spectral transformations by operating directly on downsampled raw vibration signals, reducing input size by 21 while preserving gesture-discriminative patterns. We then design two lightweight 1D convolutional architectures (1D-CNN and 1D-SepCNN) that dramatically reduce parameter counts (down to 216) compared to the 2D-CNN baseline (369 million), enabling compact yet expressive inference on resource-limited hardware. We extend the ElasticAI.Creator library to support these models for integer-only quantization and template-based RTL design generation. In addition, we integrate a hardware-aware Optuna framework to enable constraint-driven selection of model configurations that satisfy bi-objective deployment goals. Experimental results confirm that all selected configurations run under 21 ms and consume less than 3.2 mJ on the AMD Spartan-7 XC7S25 FPGA. The most efficient setting (1D-SepCNN, PS split) completes inference in under 7 ms with energy consumption smaller than 1.2 mJ, offering over 53 speedup compared to the CPU baseline [yoshida2023smatable2]. While 1D-CNN generalizes better across users and tables, 1D-SepCNN delivers superior latency and energy efficiency, highlighting a trade-off between robustness and deployment cost. Together, these results validate the feasibility of real-time gesture recognition on low-power FPGA platforms.